Can someone solve and show work for:
5w²-2w=0

Answer by sudhanshu_kmr(1152)


5w^2 - 2w =0
=> w(5w -2) = 0 [ w is common is both i.e 5w^2 and 2w]

so, w = 0
or 5w-2= 0
=> 5w = 2
=> w = 2/5
so, w = 0 or 2/5